OpenStudy (anonymous):

Given f(x) = x+3 and g(x) = X-6, where is the function f(x)/g(x) continuous?

OpenStudy (amistre64):

missing a few parts i think

OpenStudy (anonymous):

I just fixed it, sorry about that

OpenStudy (amistre64):

recall that its illegal in all 50 states and some regions of canada, to divide by zero

OpenStudy (amistre64):

this thing is NOT continuous at g(x) = 0

OpenStudy (anonymous):

Do you know how to do the formulas for each of the steps of finding the continuity of a problem?

OpenStudy (amistre64):

dunno what you mean by formulas; as is this things tells you to put the definition of g(x) under the definition of f(x); as such we get a glorified fraction that is continuous along different intervals

OpenStudy (amistre64):

g(x) = 0 when x-6=0 ; so we exclude x=6 form the domain and get: (-inf,6)U(6,inf)

OpenStudy (anonymous):

A function F is continuous at c if all three of these conditions are satisfied: 1. f(c) is defined. 2. lim f(x) = f(c) exists x->c 3. lim f(x) = f(c) x->c Those are the rules. I would have to use this to make sure the above equation is continuous.

OpenStudy (amistre64):

well, if F = f/g ; then the only hangup we have is at x=6 F(6) is not defined therefore we need not move on unless your a glutton for punishment :)

OpenStudy (anonymous):

The reason I am not getting it is because the answer says: The function f(x)/g(x) is continuous for all x except x=6. The answer is undefined (because of the zero denominator) and not continuous. So why does the answer say continuous except 6?

OpenStudy (amistre64):

... can we divide by zero? is that a mathematical do-ability?

OpenStudy (anonymous):

I have another problem that says state whether continuous at the indicated point. If its not continuous tell why. g(x) = x^2-9/x-3; x=3

OpenStudy (anonymous):

the answer for that one is not continuous, g(3) does not exist

OpenStudy (amistre64):

\[\frac{f(x)}{g(x)}=\frac{x+3}{x-6}\]

OpenStudy (amistre64):

tell me, what is 4/0 equal?

OpenStudy (anonymous):

yes i understand that.

OpenStudy (anonymous):

g(3) = 3^2-9/3-3 also makes it zero. so why is this problem not continuous and the other one is?

OpenStudy (amistre64):

are you trying to equate g(x) from one example to g(x) in another example? g(x) is just a name that gets recycled over and over again.

OpenStudy (amistre64):

you wrote: "The reason I am not getting it is because the answer says: The function f(x)/g(x) is continuous for all x except x=6." Is continuous at all EXCEPT 6, therefore it is NOT continuous at x=6 "The answer is undefined (because of the zero denominator) and not continuous. So why does the answer say continuous except 6?" EXCEPT at 6 is what you write, whch means this is NOT continuous at x=6

OpenStudy (anonymous):

LOL ok thank you. :)

OpenStudy (amistre64):

your just reading it alittle off :) happens to the best of us at times lol

OpenStudy (anonymous):

yes, plus I have been studying this for a few days. I haven't done math in 10 years and now they are making me do calculus lol. Thank you again! :)

OpenStudy (amistre64):

youre welcome, I was outta highschool 20 years before I went to college

OpenStudy (anonymous):

WOW, proud of you! keep it up! :)